# Learning Balanced Field Summaries of the Large-Scale Structure with the Neural Field Scattering Transform

### v2

@@ Abstract
We present a cosmology analysis of weak lensing convergence maps using the Neural Field Scattering Transform (NFST) to constrain cosmological parameters. The NFST extends the Wavelet Scattering Transform (WST) by incorporating trainable neural field filters while preserving rotational and translational symmetries. This setup balances flexibility with robustness, ideal for learning in limited training data regimes. We apply the NFST to 500 simulations from the CosmoGrid suite, each providing a total of 1000 square degrees of noiseless weak lensing convergence maps. We use the resulting learned field compression to model the posterior over $\Omega_{m}$, $\sigma_{8}$, and $w$ in a $w$CDM cosmology. The NFST consistently outperforms the WST benchmark, achieving a $16\%$ increase in the average posterior probability density assigned to test data. Further, the NFST improves direct parameter prediction precision on $\sigma_{8}$ by $6\%$ and $w$ by $11\%$. We also introduce a new visualization technique to interpret the learned filters in physical space and show that the NFST adapts its feature extraction to capture task-specific information. These results establish the NFST as a promising tool for extracting maximal cosmological information from the non-Gaussian information in upcoming large-scale structure surveys, without requiring large simulated training datasets.
@@ I Introduction
Extracting cosmological information from the large-scale structure requires effective compression of the various uninformative degrees of freedom in the field. For this compression, cosmology inference methods often use summary statistics: compact sets of statistics that capture selected information in the field. Traditional cosmology inference approaches develop analytic models that make predictions for how these statistics depend on cosmology, and compare these predictions to observations. The leading approach in modern large-scale structure survey analysis involves using two-point statistics, which are widely used across galaxy clustering (Masot et al., 2025), weak lensing (Secco et al., 2022; Asgari et al., 2021; Hamana et al., 2020), and joint clustering-lensing analyses (Abbott et al., 2022; Heymans et al., 2021; More et al., 2023).
@@ II.1 Data
To reflect realistic constraints on simulation resources, we intentionally conduct our analysis in a limited data regime. In practical cosmology applications, high-fidelity simulations capture more the rich non-Gaussian structure more accurately. However, increasing the fidelity of simulations often requires reducing the total number of realizations due to the heavy computational cost of accurately simulating small scale structure. Therefore, we aim to identify summary statistics that remain effective when trained on a small number of simulations. For this analysis, we choose a subset of 500 cosmologies for training (400) and validation (100), selected at random. The remaining 2000 are used as a test set to benchmark each summary statistic’s performance.
@@
Although CosmoGrid provides whole-sky convergence maps for each cosmology, we choose to use 10 of the $100~{}\mathrm{deg^{2}}$ patches, to approximate a $\sim 1000~{}\mathrm{deg^{2}}$ photometric survey such as the Kilo-Degree Survey (KiDS, Kuijken et al. (2019)). Increasing area (a possibility with future surveys such as Euclid (Euclid Collaboration et al., 2024) and LSST (Ivezić et al., 2019)) would have the effect of reducing the noise on each summary statistic. However, given that this data uses a perfect, systematic-free and noiseless reconstruction of the weak lensing field, the resulting constraint precision should only be interpreted for relative comparison purposes, not as an indication of the potential recovered precision in a KiDS-like survey.
@@ II.2 Summary Statistics
We test the ability of three summary statistics to extract cosmologically useful information from the weak lensing fields. Each summary statistic outputs a different set of coefficients that containing information about physical aspects of the field.
@@ II.3 Cosmology Inference
We benchmark the performance of the three summary statistic models for predicting cosmology parameters from the large-scale structure. We train each summary statistic model to predict $\Omega_{m}$, $\sigma_{8}$ and $w$ from the CosmoGrid weak lensing datasets, providing a test of a $w$CDM cosmology. The other cosmology and astrophysical parameters varied in CosmoGrid are unconstrained by the model, and are therefore automatically marginalized over.
@@
    Summary Statistic Pre-Training: Use raw lensing fields to predict cosmology parameters with the NFST and CNN models. Then, freeze models and compute explicit coefficients for each.
@@ II.3.1 Summary Statistic Pre-Training
We choose to pre-train the NFST and CNN summary statistics and freeze their trainable components so they produce explicit coefficients. This pre-training ensures an equivalent downstream analysis for all three summary statistic models. Additionally, we find that pre-training and computing coefficients generally improves the performance for both the NFST and CNN models compared to directly training. For pre-training, we connect each model’s output coefficients to a basic Multi-Layer Perceptron (MLP) neural network with an output for each cosmology parameter. We then train the models to predict cosmology by minimizing the mean-squared error loss between predictions and target cosmology parameters. The MLP model consists of two linear layers with 512 nodes and ReLU activation, and a final linear layer feeding to an unconstrained, 3-dimensional output corresponding to the cosmology parameters. We pre-train for 300 and 500 epochs for the NFST and CNN models respectively, with an Adam optimizer using a learning rate of $10^{-4}$. These epochs are sufficient to achieve convergence, as determined from the train and validation loss curves. For each, we use early stopping to ensure generalizable solutions. After pre-training, we freeze the trainable components in each model and apply the models to each weak lensing field, computing their corresponding set of 64 coefficients each.
@@ II.3.2 Direct Parameter Prediction
We next train a regression model to predict the respective cosmology parameters from each of the summary statistics’ coefficient sets, again by minimizing the mean-squared error loss. We use the same MLP model and training setup as used in pre-training. but train for 2000 with early stopping.
@@ II.3.3 Posterior Modeling
While predicting cosmology parameters directly is a useful indicator of summary statistic constraining power, it fails to capture the uncertainty in predictions and any degeneracies between parameters. To capture these, we instead use Neural Posterior Estimation. NPE is an approach that where a neural models learns to approximate the posterior distribution of parameters given observed data, without requiring explicit likelihood evaluation. NPE trains a neural density estimator to directly model $P(\theta|x)$ by learning from pairs of simulated parameters $\theta$ and corresponding observations $x$. It works by training a conditional density estimator to learn the mapping from data to posterior distributions. Once trained, we can sample from the NPE model to provide full posterior estimates for new observations, capturing both parameter uncertainties and degeneracies that point estimates cannot reveal (Cranmer et al., 2020). As observations $x$ we use the predicted cosmology parameters from the previous direct parameter estimation. Using these highly-compressed parameter predictions removes the compression task from the density model, which we find results in more accurate posteriors than using the raw coefficients directly.
To fit the NPE model, we use the sbi package [^1] https://sbi-dev.github.io/sbi/ and opt to use a Masked Autoregressive Flow (MAF, Papamakarios et al. (2018)) as a density model. For this setup, we perform a complete grid search over hidden_features from [16, 32, 64, 128] and num_transforms from [2, 4, 6, 8, 10, 12], finding that hidden_features=64 and num_transforms=8 produce the most accurate posterior. We emphasize that we are interested in the relative performance between the models under different input coefficients, so optimizing the posterior model further is secondary to ensuring consistency between tests. Within sbi, we use an uninformative flat prior over $[-5,5]$ for each of the normalized parameters. CosmoGrid’s finite cosmology parameter space sampling imposes an effective prior, but this is equivalent across all summary statistics, and all test cosmologies fall within this prior.
@@ III.2 Insights from NFST Filter Visualization
Figure 2: The first angular filter of the WST (top row) and trained NFST (middle row) shown in Fourier space (with coordinates $k_{x}$ and $k_{y}$), across the three scales used in the cosmology prediction. The middle of each image is the origin, with frequency increasing outwards. In the WST and NFST rows, brighter color indicates higher filter values. In the difference plot, red regions indicate an increase in the NFST’s filters relative to the WST, while blue indicates a decrease. The decreased resolution of Scale 3 is a result of filter truncation for efficiency.
@@ III.3.1 Summary Statistic-Based Field Maximization
Additionally, by supplying a specific reference field and setting $I_{C}$ such that all indices are constrained is approximately equivalent to the method described in Cheng and Ménard (2021b), providing a visualization of how a given summary statistic capture the information in a specific field.
@@ IV Discussion
The learned filters’ broader frequency pass bands suggest that cosmological information at small and intermediate scales is better extracted through statistical averaging across wide angular bands rather than precise spatial localization. This is a strong indication that fine angular resolution should not be prioritizes in wavelet-based cosmology inference, instead referencing the decrease in statistical variance that comes through averaging over more modes.
@@ V Conclusion
We demonstrate the use of the Neural Field Scattering Transform (NFST) as a summary statistic for the large-scale structure. The NFST’s balance between flexibility and robustness lets it achieve more accurate and precise cosmology constraints for cosmological inference compared to previous methods. Through trainable filters, the NFST provides an improvement in informative capacity relative to a standard Wavelet Scattering Transform (WST) model. For direct parameter prediction, the NFST provides a $6\%$ improvement in prediction RMSE for $\sigma_{8}$ and an $11\%$ improvement for $w$. For posterior modeling, the NFST provides a 17% improvement in posterior quality.
@@
We also show the direct interpretability of the NFST by comparing it to the WST at the filter-level. We see that for lensing field, Fourier-space filters with a broader angular spread are more effective for constraining cosmology. Physically, this suggests that for cosmology inference, directional information is less important than aggregation for reducing noise.
@@
Overall, these findings demonstrate that flexible and robust models such as the NFST, combined with deep learning interpretability techniques, can provide more improve the cosmological analysis process and deepen our understanding the large-scale structure.
@@ VI Acknowledgements
M.C., R.R., T.M.D., acknowledge the support of an Australian Research Council Australian Laureate Fellowship (FL180100168) funded by the Australian Government. Y.S.T is supported by the National Science Foundation under Grant No. AST-2406729. RR acknowledges financial support from the Australian Research Council through DECRA Fellowship DE240100816. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231.

### v3

@@ Abstract
We present a cosmology analysis of simulated weak lensing convergence maps using the Neural Field Scattering Transform (NFST) to constrain cosmological parameters. The NFST extends the Wavelet Scattering Transform (WST) by incorporating trainable neural field filters while preserving rotational and translational symmetries. This setup balances flexibility with robustness, ideal for learning in limited training data regimes. We apply the NFST to 500 simulations from the CosmoGrid suite, each providing a total of 1000 square degrees of noiseless weak lensing convergence maps. We use the resulting learned field compression to model the posterior over $\Omega_{m}$, $\sigma_{8}$, and $w$ in a $w$CDM cosmology. The NFST consistently outperforms the WST benchmark, achieving a $16\%$ increase in the average posterior probability density assigned to test data. Further, the NFST improves direct parameter prediction precision on $\sigma_{8}$ by $6\%$ and $w$ by $11\%$. We also introduce a new visualization technique to interpret the learned filters in physical space and show that the NFST adapts its feature extraction to capture task-specific information. These results establish the NFST as a promising tool for extracting maximal cosmological information from the non-Gaussian information in upcoming large-scale structure surveys, without requiring large simulated training datasets.
@@ I Introduction
Extracting cosmological information from the large-scale structure requires effective compression of the various uninformative degrees of freedom in the field. For this compression, cosmology inference methods often use summary statistics: compact sets of statistics that capture selected information from the field. Traditional cosmology inference approaches develop analytic models that make predictions for how these statistics depend on cosmology, and compare these predictions to observations. The leading approach in modern large-scale structure survey analysis involves using two-point statistics, which are widely used across galaxy clustering (Masot et al., 2025), weak lensing (Secco et al., 2022; Asgari et al., 2021; Hamana et al., 2020), and joint clustering-lensing analyses (Abbott et al., 2022; Heymans et al., 2021; More et al., 2023).
@@ II.1 Data
To reflect realistic constraints on simulation resources, we intentionally conduct our analysis in a limited data regime. In practical cosmology applications, high-fidelity simulations capture the rich non-Gaussian structure more accurately. However, increasing the fidelity of simulations often requires reducing the total number of realizations due to the heavy computational cost of accurately simulating small scale structure. Therefore, we aim to identify summary statistics that remain effective when trained on a small number of simulations. For this analysis, we choose a subset of 500 cosmologies for training (400) and validation (100), selected at random. The remaining 2000 are used as a test set to benchmark each summary statistic’s performance.
@@
Although CosmoGrid provides whole-sky convergence maps for each cosmology, we choose to use 10 of the $100~{}\mathrm{deg^{2}}$ patches, to approximate the survey size of a $\sim 1000~{}\mathrm{deg^{2}}$ photometric survey such as the Kilo-Degree Survey (KiDS, Kuijken et al. (2019)). Increasing area (a possibility with future surveys such as Euclid (Euclid Collaboration et al., 2024) and LSST (Ivezić et al., 2019)) would have the effect of reducing the noise on each summary statistic. However, given that this data uses a perfect, systematic-free and noiseless reconstruction of the weak lensing field, the resulting constraint precision should only be interpreted for relative comparison purposes, not as an indication of the potential recovered precision in a observational survey.
@@ II.2 Summary Statistics
We test the ability of three summary statistics to extract cosmologically useful information from the weak lensing fields. Each summary statistic outputs a different set of coefficients that contain information about physical aspects of the field.
@@ II.3 Cosmology Inference
We benchmark the performance of the three summary statistic models for predicting cosmology parameters from the large-scale structure. We train each summary statistic model to predict $\Omega_{m}$, $\sigma_{8}$ and $w$ from the CosmoGrid weak lensing datasets, demonstrating their ability to constrain $w$CDM cosmology. The other cosmology and astrophysical parameters varied in CosmoGrid are unconstrained by the model, and are therefore automatically marginalized over.
@@
    Summary Statistic Pre-Training: Use raw lensing fields to predict cosmology parameters with the NFST and CNN models. Then, freeze the models’ trainable parameters such that they behave deterministic summary statistics rather than trainable models, and compute the corresponding set of coefficients for each.
@@ II.3.1 Summary Statistic Pre-Training
The NFST and CNN models contain trainable parameters that are tuned throughout training, while the WST model is fixed and deterministic. To treat each summary statistic equally in the downstream analysis, we choose to pre-train the NFST and CNN summary statistics and freeze their trainable components. Each model then acts as a fixed function that produces an equally-sized set of coefficients, rather than an end-to-end trainable model.
Another advantage of this approach is the ability to use a different training setup for each phase. In the pre-training phase, we train each model for fewer epochs but at a higher learning rate to quickly learn useful feature representations. Then, in the downstream training phase, we train the final prediction model for more epochs at a lower learning rate to fine-tune the cosmological parameter predictions. We find that this two-stage approach consistently achieves higher performance with significantly lower compute time requirements compared to end-to-end training.
For pre-training, we connect each model’s output coefficients to a basic Multi-Layer Perceptron (MLP) neural network with an output for each cosmology parameter. We then train the models to predict cosmology by minimizing the mean-squared error loss between predictions and target cosmology parameters. The MLP model consists of two linear layers with 512 nodes and ReLU activation, and a final linear layer feeding to an unconstrained, 3-dimensional output corresponding to the cosmology parameters. We pre-train for 300 and 500 training epochs (iterations over the full training dataset) for the NFST and CNN models respectively, with an Adam optimizer using a learning rate of $10^{-3}$. For each model, the chosen number of training epochs is sufficient to achieve convergence, as determined from the train and validation loss curves. For each, we use early stopping to ensure generalizable solutions. After pre-training, we freeze the trainable components in each model and apply the models to each weak lensing field, computing their corresponding set of 64 coefficients each.
@@ II.3.2 Direct Parameter Prediction
We next train a regression model to predict the respective cosmology parameters from each of the summary statistics’ coefficient sets, again by minimizing the mean-squared error loss. We use the same MLP model and training setup as with pre-training, but instead train for 2000 epochs with a learning rate of $10^{-4}$.
@@ II.3.3 Posterior Modeling
While predicting cosmology parameters directly is a useful indicator of summary statistic constraining power, it fails to capture the uncertainty in predictions and any degeneracies between parameters. To capture these, we instead use Neural Posterior Estimation (NPE). NPE is an approach where a neural model learns to approximate the posterior distribution of parameters given observed data, without requiring explicit likelihood evaluation. NPE trains a neural density estimator to directly model $P(\theta|x)$ by learning from pairs of simulated parameters $\theta$ and corresponding observations $x$. It works by training a conditional density estimator to learn the mapping from data to posterior distributions. Once trained, we can sample from the NPE model to provide full posterior estimates for new observations, capturing both parameter uncertainties and degeneracies that point estimates cannot reveal (Cranmer et al., 2020). As observations $x$ we use the predicted cosmology parameters from the previous direct parameter estimation. Using these highly-compressed parameter predictions removes the compression task from the density model, which we find results in more accurate posteriors than using the raw coefficients directly.
To fit the NPE model, we use the sbi python package (Tejero-Cantero et al., 2020) and opt to use a Masked Autoregressive Flow (MAF, Papamakarios et al. (2018)) as a density model. For this setup, we perform a complete grid search over hidden_features from [16, 32, 64, 128] and num_transforms from [2, 4, 6, 8, 10, 12], finding that hidden_features=64 and num_transforms=8 produce the most accurate posterior. We emphasize that we are interested in the relative performance between the models under different input coefficients, so optimizing the posterior model further is secondary to ensuring consistency between tests. Within sbi, we use an uninformative flat prior over $[-5,5]$ for each of the normalized parameters. CosmoGrid’s finite cosmology parameter space sampling imposes an effective prior, but this is equivalent across all summary statistics, and all test cosmologies fall within this prior.
@@ III.2 Insights from NFST Filter Visualization
Figure 2: The first angular filter of the WST (top row) and trained NFST (middle row) shown in Fourier space (with coordinates $k_{x}$ and $k_{y}$), across the three scales used in the cosmology prediction. The middle of each image is the origin, with frequency increasing outwards. In the WST and NFST rows, brighter color indicates higher filter values. In the difference plot, red regions indicate an increase in the NFST’s filters relative to the WST, while blue indicates a decrease. Scale 2 and Scale 3 appear visually similar despite representing different scales because the Scale 3 filter’s high frequency modes are truncated for computational efficiency.
@@ III.3.1 Summary Statistic-Based Field Maximization
Additionally, supplying a specific reference field and setting $I_{C}$ such that all indices are constrained is approximately equivalent to the method described in Cheng and Ménard (2021b), providing a visualization of how a given summary statistic captures the information in a specific field.
@@ IV Discussion
The learned filters’ broader frequency pass bands suggest that cosmological information at small and intermediate scales is better extracted through statistical averaging across wide angular bands rather than precise spatial localization. This is a strong indication that fine angular resolution should not be prioritized in wavelet-based cosmology inference, instead referencing the decrease in statistical variance that comes through averaging over more modes.
@@ V Conclusion
We demonstrate the use of the Neural Field Scattering Transform (NFST) as a summary statistic for large-scale structure. The NFST’s balance between flexibility and robustness lets it achieve more accurate and precise cosmology constraints for cosmological inference compared to previous methods. Through trainable filters, the NFST provides an improvement in informative capacity relative to a standard Wavelet Scattering Transform (WST) model. For direct parameter prediction, the NFST provides a $6\%$ improvement in prediction RMSE for $\sigma_{8}$ and an $11\%$ improvement for $w$. For posterior modeling, the NFST provides a 17% improvement in posterior quality.
@@
We also show the direct interpretability of the NFST by comparing it to the WST at the filter-level. We see that for the lensing field, Fourier-space filters with a broader angular spread are more effective for constraining cosmology. Physically, this suggests that for cosmology inference, directional information is less important than aggregation for reducing noise.
@@
Overall, these findings demonstrate that flexible and robust models such as the NFST, combined with deep learning interpretability techniques, can improve the cosmological analysis process and deepen our understanding of large-scale structure.
@@ VI Acknowledgements
M.C. acknowledges the support of an Australian Government Research Training Program (RTP) Scholarship. M.C., R.R., T.M.D., acknowledge the support of an Australian Research Council Australian Laureate Fellowship (FL180100168) funded by the Australian Government. Y.S.T is supported by the National Science Foundation under Grant No. AST-2406729. R.R. acknowledges financial support from the Australian Research Council through DECRA Fellowship DE240100816. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231.